Calculate the molarity of each of the following solutions:

(a) 33.5 g of ethanol (C2H5OH) in 4.30 × 102 mL of solution

?M

(b) 22.6 g of sucrose (C12H22O11) in 78.5 mL of solution

?M

(c) 6.90 g of sodium chloride (NaCl) in 82.4 mL of solution

?M
Answer #1

Molarity is expressed as the number of moles of solute per volume of the solution. Molarity of the solutions given are calculated as follows

a. Molar mass of Ethanol C2H5OH= 12x2 + 1X5 + 16 + 1= 46g/mol

Volume= ml in litre= 4.30X102/1000

= .43L

No. Of moles = 33.5/46= .72

M= n/V

= .72/.43 = 1.69mol/L

b. Molar mass of sucrose = 342.3g/mol

Volume= 78.5/1000 = .0785L

No. Of moles = 22.6/342.3 = .066

Molarity= .066/.0785= .840mol/L

C. Molar mass of NaCl= 58.44g

Volume= 82.4/1000

=.0824L

No. Of moles = 6.90/58.44= 0.11

Molarity = n/v

= .11/.0824= 1.33mol/L.